OpenStudy (anonymous):

Find a formula for the inverse of the function and the domain of the inverse function f(x)=4+sqrt(2+11x).

OpenStudy (anonymous):

i found the inverse to be 1/11(x^2-8x+14). but the domain wont work for me. my domain was (-infinity,infinity)

OpenStudy (anonymous):

domain of \(f(x)=4+\sqrt{2+11x}\) is found by solving \[2+11x\ge0\] range is \(y\geq 4\) since \(\sqrt{2+11x}\geq 0\)

OpenStudy (anonymous):

domain of \(f^{-1}\) is the range of \(f\)

OpenStudy (anonymous):

thanks